### v1

# Graph Linearization Methods for Reasoning on Graphs with Large Language Models

Abstract
Large language models have evolved to process multiple modalities beyond text, such as images and audio, which motivates us to explore how to effectively leverage them for graph machine learning tasks. The key question, therefore, is how to transform graphs into linear sequences of tokens—a process we term “graph linearization”—so that LLMs can handle graphs naturally. We consider that graphs should be linearized meaningfully to reflect certain properties of natural language text, such as local dependency and global alignment, in order to ease contemporary LLMs, trained on trillions of textual tokens, better understand graphs. To achieve this, we developed several graph linearization methods based on graph centrality, degeneracy, and node relabeling schemes. We then investigated their effect on LLM performance in graph reasoning tasks. Experimental results on synthetic graphs demonstrate the effectiveness of our methods compared to random linearization baselines. Our work introduces novel graph representations suitable for LLMs, contributing to the potential integration of graph machine learning with the trend of multi-modal processing using a unified transformer model.
Graph Linearization Methods for Reasoning on Graphs with Large Language Models
Christos Xypolopoulos1,3, Guokan Shang2, Xiao Fei1, Giannis Nikolentzos4, Hadi Abdine2, Iakovos Evdaimon1, Michail Chatzianastasis1, Giorgos Stamou3, Michalis Vazirgiannis1,2 1Ecole Polytechnique, 2MBZUAI, 3NTUA, 4University of Peloponnese
$\dagger$\$$\dagger$\$footnotetext: These authors contributed equally.
1 Introduction
Transformer-based large pre-trained models have revolutionized machine learning research, demonstrating unprecedented performance across diverse data modalities and even a mixture of modalities, including image, audio, and text domains (Xu et al., 2023; Yin et al., 2023). Particularly, Large Language Models (LLMs) have shown promising results on arithmetic, symbolic, and logical reasoning tasks (Hendrycks et al., 2020). Despite their success, the adaptation for processing graphs—an ubiquitous data structure that encapsulates rich structural and relational information—remains a comparably emerging and underdeveloped research direction, even if it has recently been gaining attention (Ye et al., 2023; Fatemi et al., 2023; Wang et al., 2024). This asymmetry is due in large part to the inherent challenge of representing graphs as sequential tokens, in a manner conducive to the language modeling objectives typical of transformers, a challenge not encountered when dealing with the other modalities. This unique problem has encouraged us to investigate a key question: How can we represent graphs as linear sequences of tokens for transformers in a suitable way? We refer to this research endeavor as Graph Linearization.
Existing methods of using LLMs for graph machine learning tasks, such as graph reasoning and graph generation, represent entire graphs as either raw edge lists or natural language descriptions that adhere to adjacency matrices without any special treatment (Fatemi et al., 2023; Wang et al., 2024; Yao et al., 2024). For example, a natural language description of the star graph $S_{3}$ might be: “An undirected graph with nodes $a$, $b$, $c$, and $d$. Node $b$ is connected to $a$. Node $c$ is connected to $b$. Node $b$ is connected to $d$.”, its equivalent edge list representation is: “[($b$, $a$), ($c$, $b$), ($b$, $d$)]”. Either of the linearized representations is then appended with a task-specific question to form an LLM query prompt, e.g., “Is there a cycle in this graph?”. Other studies focus solely on node-level tasks (Zhao et al., 2023; Ye et al., 2023), centering the linearization around an ego-subgraph for a target node (Hamilton et al., 2017), where the neighboring graph structure and node features up to $k$-hop are described in the prompt. However, there is a lack of studies investigating how to maintain the integrity of graph structures while efficiently transforming them into sequences suitable for LLMs.
Figure 1: An example of different Graph Linearization methods applied to the motifs’ shape classification task.
Our research addresses this limitation. By relying on edge list representations as exemplified above, we study the performance impact on LLMs of various methods for ordering the edges in the list and renaming interchangeable node labels, as shown in Figure 1. We argue that the linearization of graphs should be conducted in a meaningful way, capturing properties similar to those found in natural language such as local dependency and global alignment, will benefit contemporary LLMs by enhancing their ability to understand graphs, as they are trained on trillions of textual tokens. We define local dependency as the ability to predict the next (missing) token based on the previous (surrounding) context, as seen when a sequence of tokens naturally progresses in text, analogous to the distributional hypothesis of language (Joos, 1950; Harris, 1954; Firth, 1957). Global alignment involves starting sequences from tokens with similar characteristics, aligning the sequences to mimic how text samples begin or end with common words like “The” or “In conclusion” respectively. Addressing this question helps identify LLM-suitable representations for graphs that potentially align with natural languages, unlocking new insights and applications in fields where graphs naturally represent data. Graph linearization paves the way for extending transformer models’ capabilities to graph data and unifying various graph tasks across domains, serving as a fundamental step toward building successful large graph models. This also facilitates the integration of graph learning with the multi-modal processing trend and the development of cohesive AI systems using a unified transformer model.
In this work, we developed multiple general graph linearization methods that leverage graph centrality, degeneracy, and node relabeling techniques to achieve the aforementioned natural language properties. We conducted a series of inference experiments on three basic graph reasoning tasks: node counting, max degree calculation, and motifs’ shape classification, which form the foundation for a deeper understanding of graph structures. Experimental results using the Llama 3 Instruct (Dubey et al., 2024) model on the GraphWave (Donnat et al., 2018) synthetic dataset demonstrate the effectiveness of our methods compared to random linearization baseline.
2 Related work
In this section, we first introduce previous efforts to use transformers for graph machine learning tasks. We then discuss the current trend of using modern LLMs for graph reasoning and generation. Finally, we explore linearization methods for graphs especially from specific domains.
2.1 Transformers for graph machine learning
Transformers (Vaswani et al., 2017) have been successfully applied in various domains beyond text, demonstrating both versatility and effectiveness (Devlin et al., 2018). For instance, the Vision Transformer (Dosovitskiy et al., 2020) has achieved remarkable performance in image classification tasks by treating images as sequences of patches, marking a shift from traditional CNN-based approaches. Similarly, transformers have been used in speech recognition. Models like the Speech Transformer (Dong et al., 2018) apply self-attention mechanisms to process audio data as sequences, outperforming traditional RNN-based methods.
These successes have spurred interest in using modified transformers to replace the de facto GNN-based approaches for graph machine learning tasks. Notably, Graphormer (Ying et al., 2021) enables transformer to effectively capture the dependencies and relationships within a graph. It achieves this by integrating node centrality encoding and attention biases that account for the spatial distance between nodes. Graph Transformer (GT) (Dwivedi and Bresson, 2020) generalizes the transformer architecture for graph representation learning. GT introduces the concept of relative positional encodings to account for the pairwise distances between nodes in a graph. This approach allows the model to learn rich node representations that capture both local and global graph structures.
In contrast to the above works, Kim et al. (2022) show that by treating all nodes and edges as independent tokens and inputting them into a standard Transformer encoder without any graph-specific modifications, notable outcomes can be achieved both theoretically and practically. Results on molecular graphs for quantum chemical property prediction show that this approach outperforms all GNN baselines and achieves competitive performance compared to graph Transformer variants. Despite not applying any structural alterations to the Transformer, this approach still requires sophisticated token-wise node and edge embeddings to explicitly represent the connectivity structure.
2.2 LLMs for graph reasoning
Following the recent success of LLMs in tasks beyond language processing (Hendrycks et al., 2020), several studies have explored the capacity of off-the-shelf LLMs for graph reasoning. While there is no clear consensus on the specific tasks, models are tested on understanding basic topological properties, such as graph size, node degree and connectivity, which form the foundation for a deeper understanding of graph structures (Zhang et al., 2023b). Using various prompting methods, these studies show that LLMs, even without fine-tuning, demonstrate preliminary graph reasoning abilities.
Several studies have evaluated LLMs for graph reasoning at both the node and graph levels. For example, NLGraph (Wang et al., 2024) covers eight graph reasoning tasks of varying complexity, ranging from simple tasks like connectivity and shortest path to complex problems like maximum flow and simulating graph neural networks. This work also proposes two graph-specific prompting methods that achieve notable performance improvements. GraphQA (Fatemi et al., 2023) focuses on relatively simple tasks to measure the performance of pre-trained LLMs in edge existence, node degree, node count, edge count, connected nodes, and cycle checks. It shows that larger models generally perform better on graph reasoning tasks, with graphs generated synthetically using various graph generators. Chen et al. (2023) focuses on node classification on text-attributed graphs. For a target node, this work first uses LLMs to summarize its neighborhood’s textual descriptions, which are then concatenated with the description of the target node to perform classification. Similar works include various studies that explore graph reasoning using different LLMs, prompting techniques, graph tasks, domains, and evaluation approaches (Guo et al., 2023; Zhang et al., 2023a; Hu et al., 2023; Huang et al., 2024; Liu and Wu, 2023; Das et al., 2023; Yuan et al., 2024; Wu et al., 2024; Skianis et al., 2024).
Another line of research criticizes the above approach, arguing that solely using prompt engineering or in-context learning with frozen LLMs hinders achieving top performance in downstream graph tasks. Therefore, instruction-tuning or fine-tuning is necessary. The work of Ye et al. (2023) preliminarily confirms this on the multi-class node classification task. A prompt template is designed to describe both the neighbor graph structure and node features centered around a target node up to the 3-hop level. This template, along with the correct classification answer, is used to fine-tune both encoder-decoder and decoder-only LLMs. Similarly, Zhao et al. (2023) draw inspiration from linguistic syntax trees. For a target node, the work converts its ego-subgraph into a graph syntax tree with branches describing the neighborhood’s “label” and “feature”, which are then encoded as text in the prompt. Results show that instruction-tuning performs much better that in-context learning, and is on par with GNN-based models. Similar conclusions can be observed in recent studies on graph-level reasoning settings (Luo et al., 2024).
2.3 LLMs for graph generation
Graph generation involves creating graphs with specific properties, a process that holds significant real-world value in areas like drug discovery. This task is more challenging than graph reasoning.
To the best of our knowledge, (Yao et al., 2024) was the first to show the preliminary abilities of LLMs in graph generation. The work experimented with graph generation in three settings: 1) Rule-based: generating graphs of basic structure types, given rules describing the desired structures, e.g., trees, cycles, etc.; 2) Distribution-based: generating graphs following a structural type distribution $p$, given a set of example graphs with the same distribution; 3) Property-based: generating molecule structures with specific properties, given example molecules in SMILES format. Results show that LLMs have reasonably good abilities across the three tasks, and advanced prompting techniques do not necessarily lead to better performance.
2.4 Linearization for specific graphs
In deriving an ordering for graphs, topological sorting in graph theory examines the linear ordering of directed acyclic graphs, such that for every directed edge $(u,v)$, $u$ precedes $v$ in the ordering. However, such graph traversal is node-centric, making edge information not encoded.
In other domains involving specific types of graphs, such as discourse graphs—a directed weakly connected graph reflecting discourse structure—nodes represent utterances, and edges represent discourse relations (e.g., elaboration, clarification, completion) within a conversation Rennard et al. (2024). The work of Chernyavskiy et al. (2024) proposes a linearization method for discourse graphs that arranges utterances chronologically, assigning unique identifiers to speakers, utterances, and addressees. It incorporates discourse relations and sentiment tokens to generate a structured sequence, using special tokens for clarity. This structured sequence is then used to train a BART (Lewis et al., 2020) for dialogue generation. Similarly, Abstract Meaning Representation (AMR) uses directed acyclic graphs to provide a structured semantic representation of language, incorporating semantic roles with annotated arguments and values where nodes represent concepts and edges represent semantic relations (Banarescu et al., 2013). AMR corpora are usually linearized using the PENMAN-based notation (Patten, 1993) as in the work of (Ribeiro et al., 2021) and (Hoyle et al., 2021) to fine-tune pre-trained language models to perform graph-to-text generation. For citation networks, Guo et al. (2023) have explored the Graph Modelling Language (GML) and Graph Markup Language (GraphML) for graph representation (Himsolt, 1997; Brandes et al., 2013). GML is a simple, human-readable format, while GraphML is XML-based and offers extensibility for complex applications.
The scope of our work. Unlike the above linearization methods limited to specific types of graphs, where linearization can be naturally derived to some extent, we focus on general graphs. Furthermore, unlike previous works using LLMs for reasoning and generation tasks, where edge lists are directly leveraged without special treatment, we introduce various linearization methods for ordering the edges in the list and renaming interchangeable node labels to make them suitable for LLMs. Although our work does not involve experiments with graph generation or instruction tuning, our graph linearization methods are general and applicable to various scenarios. This allows for the effective transformation of graph structures into sequences suitable for language models and has the potential to improve performance in both reasoning and generation tasks, with or without fine-tuning.
3 Graph linearization methods
Generally speaking, we define graph linearization as the process of representing graphs as linear sequences of tokens. In this work, we aim to identify the linearization approaches that will benefit state-of-the-art LLMs by enhancing their ability to understand graphs. We argue that linearized graphs, represented as sequences of tokens, should capture properties similar to those in natural language, given the fact that LLMs are pre-trained on trillions of textual tokens. Such properties should include local dependency and global alignment.
Local dependency refers to the ability to predict the next (missing) token based on the previous (surrounding) context, within the token sequence of a single linearized graph. This property is analogous to the fundamental distributional hypothesis of language (Joos, 1950; Harris, 1954; Firth, 1957), which states that words that occur in similar contexts tend to have similar meanings (or functions). This hypothesis suggests that given a new word, one should be able to figure out its meaning based on the contexts in which it is used. In fact, the masked and casual language modeling for training encoder-only (Devlin et al., 2019) and decoder-only (Radford et al., 2019) language models, can be seen as instantiations of this hypothesis. Global alignment refers to the alignment across the token sequences of different linearized graphs, so they start or end with tokens sharing similar natures. For example, linearization can be always start from the node with the highest degree, and such nodes are all relabelled as index 0 for different graphs. This property is to reflect the fact that text documents often begin or end with common words like “The” or “In conclusion”.
Figure 2: $k$-core decomposition.
Figure 3: Overview of GraphWave synthetic dataset.
By relying on edge list, which defines a graph in terms of its individual connections, we study the performance impact on LLMs of various methods for ordering the edges in the list and renaming interchangeable node labels, as shown in Figure 1. More specificly, we leverage the advances in graph degeneracy and centrality as detailed in below to meet the local dependency property.
Graph degeneracy (Seidman, 1983). Let $G(V,E)$ be an undirected graph with $n=\left|{V}\right|$ nodes and $m=\left|{E}\right|$ edges. A $k$-core of $G$ is a maximal subgraph of $G$ in which every vertex $v$ has at least degree $k$. As shown in Figure 2, the $k$-core decomposition of $G$ forms a hierarchy of nested subgraphs whose cohesiveness and size respectively increase and decrease with $k$. The higher-level cores can be viewed as a filtered version of the graph that captures main information. The Core Number of a node is the highest order of a core that contains this node.
Graph centrality. A key concept in network analysis to determine the influence or importance of nodes based on the structure of the graph. Degree centrality is one of the simplest measures of node importance Freeman (1978). It is defined as the number of edges incident to a node. For an undirected graph with $n=\left|{V}\right|$ nodes and $m=\left|{E}\right|$ edges, the degree centrality $C_{D}(v)$ of a vertex $v\in V$ is calculated as $C_{D}(v)=\deg(v)=\sum_{u\in V}A_{vu}$ where $A$ is the adjacency matrix of the graph, and $A_{vu}=1$ if there is an edge between vertices $v$ and $u$, and 00 otherwise. PageRank is a more sophisticated centrality measure, originally developed by Brin and Page (1998) for ranking web pages. It extends the concept of degree centrality by considering not only the number of links a node has but also the importance of the nodes linking to it. PageRank, although designed for directed graphs, can also be adapted for undirected ones by treating all edges as bidirectional. The PageRank centrality $C_{PR}(v)$ of a node $v\in V$ is computed iteratively using the formula: $C_{PR}(v)=\frac{1-\alpha}{|V|}+\alpha\sum_{u\in\mathcal{N}(v)}\frac{C_{PR}(u)}{\deg^{+}(u)}$ where $\alpha$ is a damping factor typically set to $0.85$, $\mathcal{N}(v)$ represents the set of nodes linking to $v$, and $\deg^{+}(u)$ is the out-degree of node $u$. PageRank centrality effectively captures the notion that connections from highly-ranked nodes contribute more to the ranking of a given node than connections from low-ranked nodes Page et al. (1999).
Our linearization methods. We order the edges (node pairs) in the edge list based on Core Number, Degree and PageRank as described in above. Specifically, first we rank nodes by their corresponding property. Then, we start with the node having the highest property value and select the edges connected to it, arranging these edges in a random order. This process is repeated for nodes with the next highest property values, continuing in descending order. In the case where two or more nodes share an equal value, the order is selected randomly. After the ordering process has concluded, each edge list constitutes a sequence of tokens following a descending order of node importance. Additionally, we examined the case where we order the edges directly instead of the nodes. To do this, we transformed each graph into its corresponding Linegraph. A linegraph $L(G)$ of a graph $G$ is the graph where each edge of $G$ is replaced by a node, and where two edges of $G$ are connected in $L(G)$ if they are incident in $G$. Then we applied the process described previously using again the core number of each node in $L(G)$.
Finally, we rely on node relabeling to achieve the property of global alignment. For example, the node with index 0 constantly is the one that has the highest core number and so on. This might be beneficial for the LLM to consistently maintain a connection between node indexes and their corresponding importance properties. We evaluate all aforementioned linearization methods with and without this node relabeling strategy.
4 Dataset
We constructed a synthetic graph dataset using GraphWave (Donnat et al., 2018). This graph generator was originally developed for controlled experimentation and evaluation of node embedding techniques and measurement of structural equivalence on graphs with known network motifs. Motifs, in the context of network science, are sub-graphs that repeat themselves either within the same graph or across different graphs. For our analysis, these structurally similar sub-graphs enable us to construct structure-related tasks, that allow us to assess the ability of language models to analyze and infer structural features within these graphs.
The generator operates by constructing sequentially a base graph, that follows either a cycle or chain structure, and then attaching a number of motifs that follow predetermined shapes, to form the final graphs. From the available motif’s shapes, we chose to adopt five of them—cliques, stars, fans, diamonds, and trees—excluding shapes that share structural similarities with the selected ones. To cover a wider variety of structural complexities, we also include all the combinations of two shapes per graph, along with all combinations of three shapes with unique shapes per triplet. Example graphs can be found in Figure 3.
For each combination of shapes, we generated 100 graphs, leading to a total of 3000 graphs. This ensures sufficient variance in graph sizes and in the combinations of base and motif sub-graphs. The number of nodes in each graph shape is selected randomly, with the following constraints: base (3-21 nodes); clique, fan, and star (4-11 nodes); diamond (6 nodes); and tree (perfect binary trees with 3-6 levels). The generated dataset contains an average of 32.33 nodes and 43.72 edges per graph.
Node Counting Max Degree Calculation
Motifs’ Shape
Classification
Accuracy MAE Accuracy MAE Accuracy
Linearization
Method
ZS OS ZS OS ZS OS ZS OS ZS OS
Default Node Labels CoreNumber 43.39 45.7 1.16 1.12 19.01 20.55 1.93 2.04 39.2 38
Degree 53.85 62.28 1.03 1.12 23.67 30.89 1.50 1.70 46.1 39.57
PageRank 50.92 51.38 2.09 1.9 20.34 29.92 1.67 1.63 47.27 33.5
Linegraph 41.59 34.71 1.52 1.92 19.21 11.74 1.77 2.05 44.63 33.5
Node Relabeling CoreNumber 42.1 28.39 0.9 1.55 13.70 17.51 2.09 1.95 41.43 46.9
Degree 57.70 54.58 0.66 0.74 24.64 28.32 1.45 1.38 44.13 43.03
PageRank 47.75 42.21 1.58 1.57 18.21 19.61 1.77 1.81 46.5 35.37
Linegraph 34.34 26.18 0.97 1.14 14.30 11.27 2.04 2.15 45 26.8
Baselines MotifAware 45.92 37.48 1.15 1.35 10.40 11.01 3.74 4.18 49.37 31.93
Random 48.95 41.42 0.6 0.8 11.55 8.80 2.55 2.60 43.95 31.89
Table 1: Comparison of various linearization methods across all tasks using Llama 3 Instruct 8B. The table presents the performance of different linearization methods, including the default node ID labeling against the node relabeling version, evaluated under both zero-shot (ZS) and one-shot (OS) settings. Metrics used for comparison are mean absolute error (MAE) and accuracy, where a lower MAE indicates more precise predictions, and higher accuracy reflects the model’s ability to correctly respond to the given task. Bold indicates the best performance in each setting, while underlined scores represent the second best method in those settings.
5 Experimental setup
In this section, we describe the protocol we followed to test the extent to which our linearization methods perform on graph reasoning tasks.
Tasks and prompts. Our experiments include three graph reasoning tasks. The first task is node counting, where the LLM infers the number of nodes from a linearized graph’s edge list. The second task is max degree calculation, where the LLM determines the maximum node degree. The third task, motifs’ shape classification, gives the LLM detailed definitions of the five motifs’ shape in our dataset, and ask it to identify which motif is present in a graph. This combination of reasoning and classification tasks allows us to comprehensively evaluate LLMs’ understanding of structural features and examine how different edge list orderings affect their performance.
In every prompt, a node $i$ is represented by an incremental integer while an edge between nodes $i$ and $j$ is denoted by the bracketed pair $(i,j)$. An edge list is expressed as a sequence of edges, sorted according to the scheme used in each linearization method. We tested both zero-shot and one-shot approaches, where a randomly selected graph from the dataset was used consistently as the one-shot example across all experiments. The prompt templates are provided in Appendix A.
LLMs. For our experiments, we employed the 8B parameter version of Llama 3 Instruct (Dubey et al., 2024), quantized to 4 bits. We opted for $1\mathrm{e}{-3}$ temperature and $1\mathrm{e}{-1}$ sampling parameter values, in order to ensure more deterministic and coherent outputs, reducing randomness and enhancing consistency in generated responses. This allows us to assess its sensitivity to our linearization methods. The experiments were run on an NVIDIA A5000 GPU with 24GB of VRAM.
Baselines. For our comparisons, we used two baseline linearization methods.
MotifAware. This baseline uses the default ordering in which the graph is generated. As the graphs are built incrementally, starting with the base graph and then motifs are attached, this method produces a structure-aware edge list order (see Figure 1 “Default”). We expect this embedded structural knowledge to reflect on our tasks and increase LLM performance. However, generating such structure-aware lists requires knowledge of the graph construction process, which may not be feasible in real-world scenarios with larger, more complex graphs.
Random. This baseline involves a fully Random ordering of the edge list, where edges are arranged without following any inherent scheme. To further eliminate structural biases, we also randomly shuffle the node IDs (labels). To mitigate the risk of skewed results, we applied five different random orderings and averaged their performance.
Evaluation Metrics: We use exact accuracy to compare our methods, measuring the ratio of correct predictions as $\frac{1}{n}\sum_{i=1}^{n}I\left(y_{i}=\hat{y}_{i}\right)$, where $n$ is the number of graphs, $y_{i}$ the correct answer, and $\hat{y}_{i}$ the LLM’s response. For the counting tasks, we consider a result accurate only if it is an exact match. For the classification task, accuracy reflects the total across all shapes, requiring the predicted shape to appear at least once in the graph.
While accuracy provides a strict measure of correctness, it does not quantify the average magnitude of errors. To gain more fine-grained insights for counting tasks, especially when exact matches are not achieved, we also use Mean Absolute Error (MAE), calculated as $\frac{1}{n}\sum_{i=1}^{n}\left|y_{i}-\hat{y}_{i}\right|$, where $n$ is the number of graphs, $y_{i}$ the actual number of nodes (or max degree) and $\hat{y}_{i}$ the LLM’s response.
6 Experimental results analysis
In this section, we compare the performance across different tasks and linearization methods. The results are shown in Table 1.
Node counting. All methods show relatively low average errors. This suggests that the model is generally accurate in this task. However, models with lower average errors do not always show higher accuracy when compared. This indicates that while some methods frequently predict the correct node count, their errors are larger when they do occur.
Our proposed linearization methods consistently outperform the baselines in accuracy, with the Degree and PageRank methods showing the most potential. The Random baseline and the Relabeled Degree method exhibit nearly identical average errors, with one performing better in the one-shot setting and the other in the zero-shot setting. Although the Random baseline has a similarly low average error compared to the Relabeled Degree method, it suffers a significant drop in accuracy.
Notably, the node relabeling strategy leads to a performance drop in accuracy, except for the zero-shot Degree method. However, applying relabeling reduces the average error, except for the one-shot CoreNumber method. A consistent performance drop is observed in the one-shot setting across all methods. Finally, centrality-based methods tend to achieve higher accuracy than degeneracy-based approaches.
Max degree calculation. This task is more complex than the node counting task. Node counting only requires the model to identify the largest number in a sequence, whereas max degree calculation requires counting how often each node id appears and determining which id occurs most frequently. This additional layer of computation—counting and comparing occurrences—makes max degree calculation more challenging for LLMs. This complexity is reflected in the performance results across different setups.
With default node labeling, the Degree and PageRank methods stands out in one-shot setting, achieving the highest accuracy. The rest of the methods, while achieving competing average errors, maintain low accuracy scores.
When node relabeling is applied, the Degree method shows notable improvement in zero-shot, with increased accuracy and lower MAE. However, the other methods remain below 20% accuracy. Additionally, we observe that both baselines perform poorly, achieving very low accuracy.
These results highlight that Degree-based linearization is particularly effective for the max degree calculation task allowing it to handle the frequency-based complexity better than the others.
Motifs’ shape classification. For this task, we observe that the relabeling strategy has varied effects, leading to slightly higher accuracy scores across our methods mainly in the one-shot setting. However, similar to the node counting task, there is a consistent decline in performance from the zero-shot to the one-shot setting.
The MotifAware baseline stands out, performing exceptionally well in the zero-shot setting due to its embedded structural information, which likely enhances the model’s ability to recognize motif patterns. The PageRank method, however, performs on par. Unlike in the node counting task, centrality-based methods seem to yield accuracy comparable to that of the degeneracy-based ones.
7 Conclusion
Our research highlights the potential of large language models to extend their capabilities beyond text to encompass graph reasoning tasks. Central to this integration is the concept of “graph linearization”, where graphs are transformed into linear sequences of tokens, enabling LLMs to process them in a manner similar to natural language. We emphasize that effective graph linearization should preserve certain properties of natural language, such as local dependency and global alignment, to leverage the vast training LLMs have undergone with textual data. To this end, we developed and evaluated several graph linearization techniques grounded in graph centrality, degeneracy, and node relabeling strategies. Our experimental results on synthetic graphs underscore the superiority of these methods over random linearization approaches. This work paves the way for novel graph representations that are compatible with LLMs, fostering the integration of graph reasoning and machine learning within the broader scope of multi-modal processing in unified transformer models.
Limitations
Our study has a few limitations that future work could address. First, limited structural features were considered. We did not incorporate certain important graph features, such as communities and connected components, which may improve the method’s ability to capture complex structural patterns. Second, dataset diversity is limited. The experiments were conducted on a small set of synthetic graphs, so future research should explore a wider variety of graph datasets to ensure better generalizability. Third, task variety was narrow. While we focused on tasks like node counting and motif classification, future work could extend to more advanced tasks, such as graph clustering, matching, or even graph generation with fine-tuning.
Appendix A Prompt templates
Node Counting:
In an undirected graph, (i, j) means that node i and node j are connected with an undirected edge. The edges in G are: {linearized graph}
Q: How many nodes are in this graph?
Max Degree Calculation:
In an undirected graph, (i, j) means that node i and node j are connected with an undirected edge. The degree of a node is the number of edges connected to the node. Given a graph G and it’s list of edges respond to the following question:
Q: Without any justification, what is the maximum node degree in the following graph G? G: {linearized graph}
Motifs’ Shape Classification:
In an undirected graph, (i, j) means that node i and node j are connected with an undirected edge. The graph contains a motif graph with strictly one of the following structures. {structure}: {definition}
Q: Which of the defined structures is included in the following graph? graph: {linearized graph}